OpenStudy (anonymous):

How many solutions does 2(x-5)=2x+3 have

OpenStudy (anonymous):

There are no solutions for x.

OpenStudy (anonymous):

\[2(x-5)=2x+3\]\[2x - 10 = 2x+3\]\[2x - 2x = 3+10\]\[0\neq13\]